OpenStudy (anonymous):

I will Give out a medal and best response! . Simplify the following problem using the distributive property: 3x + 4y + 3(x – y) – 4x. Show each step in your calculation. Justify each of your steps using number properties. Answer: Work: Number Property Us

OpenStudy (anonymous):

Somebody Help me!!!! :(

OpenStudy (anonymous):

. Simplify the following problem using the distributive property: 3x + 4y + 3(x – y) – 4x. Show each step in your calculation. Justify each of your steps using number properties. show the Work: Number Property Used

OpenStudy (anonymous):

I'll have a go at solving this. The distributive property can be used on the 3(x-y) portion. If you distribute the 3 it will create 3x-3y instead of 3(x-y)

OpenStudy (anonymous):

Then you combine like terms. Do you need help with the combining like terms or can you do that?

OpenStudy (anonymous):

Could you?

OpenStudy (anonymous):

I could explain it to the best of my ability. All our X values can be combined. All the Y values can be combined. But we Can't combine X and Y. So for example the X values are: 3x 3x and -4x so our equation to solve this is 3x+3x-4x. Combined we have 2x. Can you do this for the Y values or is it still unclear?

OpenStudy (anonymous):

it's still unclear. sorry!

OpenStudy (anonymous):

That's completely okay! So our Y values are 4y and -3y. We can basically ignore our y for now and work with the numbers. It's elementary solving at this point. 4-3=1. we have 1y or just y. Your simplified equation becomes 2x+y!!!

OpenStudy (anonymous):

So 2x+y is the simplified version of this problem? Thank you! your the best!

OpenStudy (anonymous):

Correct! Glad I could help you!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

3 x + 4 y + 3 x - 3 y - 4 x = y - 4 x , 3(x-y) = 3x - 3y 3x - 3x = 0 4y - 3y = y -4 x = -4x final equation : -4x + y

OpenStudy (anonymous):

thanks @Ogzii YOUR THE BEST!!!!! ;p

OpenStudy (anonymous):

@Ogzii could you please let me know the number property used?

OpenStudy (anonymous):

@Ogzii Do you mind explaining? I don't see where the -3x comes from.

OpenStudy (anonymous):

i guess i was wrong its , y + 2x